OpenStudy (anonymous):

In the triangle below, determine the value of a. 2.5 11.7 9.6 2.3

OpenStudy (anonymous):

can someone help with the formula

OpenStudy (anonymous):

thats not one of the choices @moneymob2434

OpenStudy (anonymous):

ohh sorry lol its \[a ^2+b ^{2} =c ^{2}\]

OpenStudy (anonymous):

i cant do that because there only one value lol @moneymob2434

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

All triangles equal a sum of 180° so all you have to do is 180-15 to fin the side of 1 side I think

OpenStudy (anonymous):

its not one of the choices tho @pizza12

OpenStudy (anonymous):

Then plug it into the equation a^2 +b^2=c^2

OpenStudy (anonymous):

@Butterfield1215

OpenStudy (butterfield1215):

one sec

OpenStudy (butterfield1215):

11.7

OpenStudy (butterfield1215):

@nicoleee23

OpenStudy (anonymous):

thankyou soo much @Butterfield1215 i have a couple more

OpenStudy (butterfield1215):

ok